3.1.4. Find all separable eigensolutions to the following partial differential equations:
(b) ut = ux

(c) ut = xux.
Answer #1

© ¥= x Ubc-80 Let U14H)= X(X) TCH be solution of o Then 4 = x+4 Ux = XT Bom o x=xxt I = xx = 6(803) = I=x + XX=K JI=. U = Ux0 Let UCSIA = XT be Bolition cro - Them U2=X7 & Ux=xT Proin 0, xtext > = I= K (8) = }}kſiax & SFexflot lgx = kx